HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THE NEWLY SET-UP SPECTROSCOPIC MOLECULAR BEAM APPARATUS:

PRODUCING TRANSIENT MOLECULES IN A COLD SUPERSONIC JET

\underline{S. SCHWETJE} , \underline{J. JAKOB} , \underline{V. GRIGORIAN} , M. A. KAUFMANN, A. A.

BREIER, G. W. FUCHS and T. F. GIESEN, Laborastrophysik, Universität Kassel, Heinrich-Plett-Straße 40, 34132 Kassel, Germany

Diffuse interstellar bands, or DIBs, are features in absorption spectra of astronomical objects. Probable carriers of those bands are various, yet unassigned molecules found in the visible and near-infrared sprectra of interstellar gas and dust clouds. Complex organic molecules may contribute to the DIBs as shown in recent work by Campbell et al. when they assigned some of the near-infrared DIBs to the molecular ion C_{60}^{+}.^{1} To assign more DIBs, laboratory experiments are required. Therefore, it is necessary to employ a powerful source which is capable of producing astrochemically relevant molecules coupled with high resolution spectroscopy.

The Even-Lavie high-pressure pulsed valve provides a supersonic jet with cooling properties below 1 K, backing pressures up to 100 bars, repetition rates up to 600 Hz and gas pulse widths of 20 microseconds.^{2} Together with a mountable AC discharge source it is capable of producing both stable and transient molecules in a jet to approach the production of astrochemically relevant species.

Here we present preliminary data from a molecular beam setup coupled with a discharge source and REMPI spectroscopy as the detection technique.

1doi:10.1038/nature14566, E. Campbell, M. Holz, D. Gerlich et al., “Laboratory confirmation of C_{60}^{+} as the carrier of two diffuse interstellar bands”, Nature 523, 322-323 (2015).

2doi:10.1063/1.481405, U. Even, J. Jortner, D. Noy, N. Lavie, C. Cossart-Magos, “Cooling of large molecules below 1 K and He clusters formation”, J. Chem. Phys. 8 May 2000; 112 (18): 8068-8071.

p-number: p226  Submitted on Mon, 19 Jun 2023 11:59:28 +0200